OpenStudy (anonymous):

evaluate the integral: xe^(2x^2) dex

OpenStudy (anonymous):

I answered the other one :P

OpenStudy (anonymous):

4x^2=u du =8xdx integral1/8e^u=1/8e^(4x^2)+c